# Cosmology on the Generalized Proca Theory

### v1

@@ 2.2.2 Dynamical System
which is the usual form of the continuity equation of an uncoupled fluid, as it was described on the Section 1.1.2, stressing that we are considering that DE can be an anisotropic fluid.
@@ 2.2.3 Analytical Fixed points
where we have replaced the expression for $q$ in terms of the dynamical variables given in Eq. (2.10), and some algebraic manipulations have been performed. Notice that these equations compose a polynomial system of degree greater than $4$, since Eq. (2.25) is an equation of degree $6$ in the variable $\Sigma$. This means that the system is not analytically solvable in general due to the fundamental theorem of Galois theory [71]. However, in the specific case where $\Sigma=0$, Eq. (2.25) reduces to $z=0$, and Eq. (2.24) is trivially satisfied. This is consistent with the fact that the vector field is the source of the anisotropy, as can be seen in Eq. (2.6).
Under this simplification, the reduced system, consisting of Eqs. (2.22), (2.23), and (2.26), becomes a polynomial system of degree $4$. In such cases, analytical solutions exist, as we previously demonstrated for the tachyon field in the FLRW metric in Section 1.7. Therefore, we expect to obtain analogous results under this simplification.
@@ 2.3.1 Isotropic Dark Energy Attractor
$\alpha=0.5,\quad\beta=0.1,$ (2.41)
@@
In Figure 2.5, we plot the evolution of the density parameters, the effective equation of state of the universe, and the equation of state of dark energy (DE) as functions of the redshift $z_{r}$. The early universe ($z_{r}>10^{6}$) is dominated by radiation, as indicated by the red dotted line. At $z_{r}\approx 3200$, the radiation-matter equality occurs, where $\Omega_{m}\simeq\Omega_{r}$. Following this transition, the DE contribution is $\Omega_{\text{DE}}\approx 9.89\times 10^{-10}$, which satisfies the Big Bang Nucleosynthesis (BBN) constraint, $\Omega_{\text{DE}}<0.045$ at $z_{r}=1200$ [97]. From this transition until $z_{r}\approx 0.3$, the universe is matter-dominated, as shown by the light brown dashed line. At $z_{r}=50$, the DE contribution to the energy budget is $\Omega_{\text{DE}}=1.74\times 10^{-5}$, consistent with the Cosmic Microwave Background (CMB) constraint, $\Omega_{\text{DE}}<0.02$ at this redshift [1]. Subsequently, the Universe transitions to a DE-dominated phase, represented by the black solid line, during which the expansion accelerates as $w_{\text{eff}}<-1/3$ (blue dot-dashed line). The behavior of the DE equation of state matches theoretical expectations, as discussed in Section 1.7.4: it remains $w_{\text{DE}}=-1$ during the radiation and matter-dominated epochs but takes on a different value during the DE-dominated era. This value depends solely on $\alpha$ and, in this case, is given by $w_{\text{eff}}=w_{\text{DE}}\simeq-0.92$. Additionally, in Figure 2.6, we plot $w_{\text{DE}}$ for several values of $\alpha$ while keeping $\beta$ fixed, demonstrating that $w_{\text{DE}}\rightarrow-1$ as $\alpha\rightarrow 0$, consistent with theoretical predictions.
@@ 2.4 Conclusions
In this chapter, we have put forward a numerical method to explore the parameter space of a cosmological model when no analytical fixed points are available. We applied our method to a specific model of anisotropic dark energy based on the interaction between a scalar tachyon field and a vector field in a Bianchi I background. We have explicitly shown that the anisotropic attractor of the system has no analytical description, given that the degree of the algebraic system from which this point must be computed is greater than four [see Eqs. (2.22)-(2.25)]. However, when the anisotropy is neglected, the system is reduced to a system of degree 4 and thus analytical solutions exist, which we presented in Eqs. (2.27), (2.29) and (2.31), as well as in the Section 1.7. In particular, our method allowed us to find the parameter space of the model where anisotropic accelerated solutions exist as attractors of the system, which we plot in Figure 2.4. Then, we checked the consistency of the method by numerically solving the full autonomous system in Eqs. (2.15)-(2.19) for a particular set of initial conditions. As a last remark, we would like to stress on the generality of our method, as explained in Section 1.6.1, i.e., in principle, it can be applied to any DE scenario.
@@ 3.1 Introduction
In Chapter 1, we presented a concise overview of cosmological studies, emphasizing the $\Lambda$CDM model, illustrating the application of a simple dynamical systems approach to a toy model, and providing a brief introduction to the Generalized SU(2) Proca theory. We demonstrated how even minor modifications to a model can significantly increase its complexity, necessitating advanced methods to derive meaningful insights and reconstruct the cosmic history. Chapter 2 further developed these concepts by focusing on an anisotropic Tachyon field, showcasing the use of numerical techniques within the dynamical systems framework to analyze its behavior and explore its implications for cosmic evolution.
@@ 3.2.3 Stability Analysis
As discussed in Section 1.7.3, the stability of fixed points can be determined by analyzing the eigenvalues of the Jacobian matrix, defined as: [^4] Alternative techniques, such as Lyapunov exponents [43] or Malkin’s criterion [133], may be required if linear stability analysis fails.
@@ 3.4 Conclusions
In this thesis, we employed the dynamical systems approach to explore two different cosmological models, each addressing distinct challenges in understanding the expansion history of the Universe. In Chapter 2, we developed a numerical framework for analyzing anisotropic dark energy models, specifically focusing on the interaction between a scalar tachyon field and a vector field in a Bianchi I background. This method, which does not rely on analytical fixed points, enabled the identification of regions within the parameter space where accelerated solutions could act as attractors of the system. The flexibility of this approach makes it applicable to any dark energy scenario, thereby demonstrating its general applicability.
In Chapter 3, we investigated the Generalized SU(2) Proca (GSU2P) theory, with particular focus on its potential to drive both primordial inflation and late-time cosmic acceleration. While the dynamical systems approach revealed regions where accelerated solutions could theoretically exist, it ultimately helped us rule out the GSU2P model as a viable candidate for a complete cosmological acceleration history. The model’s failure to produce stable attractors and maintain consistent cosmological evolution across different epochs prevented a smooth transition from inflation to a radiation-dominated phase, as well as from a matter-dominated phase to accelerated expansion. Thus, the dynamical systems approach served to discard this model as a viable candidate for cosmological history, highlighting its limitations in replicating the observed evolution of the Universe. These results emphasize the importance of refining theoretical models to align with observational data while addressing potential instabilities that may arise in their formulation. In particular, the presence of singularities in phase space, such as those resulting from the regularization of dynamical equations, can disrupt the integrability and evolution of cosmological models. Depending on the behavior of the system during numerical integration, such singularities can halt its progression or introduce additional instabilities, ultimately limiting the model’s ability to replicate a consistent and observationally accurate cosmological history. These challenges underscore the need for rigorous examination and refinement of models to ensure their robustness across different epochs of cosmic evolution.

### v2

@@ 2.2.2 Dynamical System
which is the usual form of the continuity equation of an uncoupled fluid, as it was described on the Section 1.2, stressing that we are considering that DE can be an anisotropic fluid.
@@ 2.2.3 Analytical Fixed points
where we have replaced the expression for $q$ in terms of the dynamical variables given in Eq. (2.10), and some algebraic manipulations have been performed. Notice that these equations compose a polynomial system of degree greater than $4$, since Eq. (2.25) is an equation of degree $6$ in the variable $\Sigma$. This means that the system is not analytically solvable in general due to the fundamental theorem of Galois theory [70]. However, in the specific case where $\Sigma=0$, Eq. (2.25) reduces to $z=0$, and Eq. (2.24) is trivially satisfied. This is consistent with the fact that the vector field is the source of the anisotropy, as can be seen in Eq. (2.6). Under this simplification, the reduced system, consisting of Eqs. (2.22), (2.23), and (2.26), becomes a polynomial system of degree $4$. In such cases, analytical solutions exist.
@@ 2.3.1 Isotropic Dark Energy Attractor
$\alpha=0.5,\quad\beta=0.1.$ (2.41)
@@
In Figure 2.5, we plot the evolution of the density parameters, the effective equation of state of the universe, and the equation of state of dark energy (DE) as functions of the redshift $z_{r}$. The early universe ($z_{r}>10^{6}$) is dominated by radiation, as indicated by the red dotted line. At $z_{r}\approx 3200$, the radiation-matter equality occurs, where $\Omega_{m}\simeq\Omega_{r}$. Following this transition, the DE contribution is $\Omega_{\text{DE}}\approx 9.89\times 10^{-10}$, which satisfies the Big Bang Nucleosynthesis (BBN) constraint, $\Omega_{\text{DE}}<0.045$ at $z_{r}=1200$ [101]. From this transition until $z_{r}\approx 0.3$, the universe is matter-dominated, as shown by the light brown dashed line. At $z_{r}=50$, the DE contribution to the energy budget is $\Omega_{\text{DE}}=1.74\times 10^{-5}$, consistent with the Cosmic Microwave Background (CMB) constraint, $\Omega_{\text{DE}}<0.02$ at this redshift [7]. Subsequently, the Universe transitions to a DE-dominated phase, represented by the black solid line, during which the expansion accelerates as $w_{\text{eff}}<-1/3$ (blue dot-dashed line). The behavior of the DE equation of state matches theoretical expectations: it remains $w_{\text{DE}}=-1$ during the radiation and matter-dominated epochs but takes on a different value during the DE-dominated era. This value depends solely on $\alpha$ and, in this case, is given by $w_{\text{eff}}=w_{\text{DE}}\simeq-0.92$. Additionally, in Figure 2.6, we plot $w_{\text{DE}}$ for several values of $\alpha$ while keeping $\beta$ fixed, demonstrating that $w_{\text{DE}}\rightarrow-1$ as $\alpha\rightarrow 0$, consistent with theoretical predictions.
@@ 2.4 Conclusions
In this chapter, we have put forward a numerical method to explore the parameter space of a cosmological model when no analytical fixed points are available. We applied our method to a specific model of anisotropic dark energy based on the interaction between a scalar tachyon field and a vector field in a Bianchi I background. We have explicitly shown that the anisotropic attractor of the system has no analytical description, given that the degree of the algebraic system from which this point must be computed is greater than four [see Eqs. (2.22)-(2.25)]. However, when the anisotropy is neglected, the system is reduced to a system of degree 4 and thus analytical solutions exist, which we presented in Eqs. (2.27), (2.29) and (2.31). In particular, our method allowed us to find the parameter space of the model where anisotropic accelerated solutions exist as attractors of the system, which we plot in Figure 2.4. Then, we checked the consistency of the method by numerically solving the full autonomous system in Eqs. (2.15)-(2.19) for a particular set of initial conditions. As a last remark, we would like to stress on the generality of our method, as explained in Section 1.5.1, i.e., in principle, it can be applied to any DE scenario.
@@ 3.1 Introduction
In Chapter 1, we presented an overview of cosmological studies, emphasizing the $\Lambda$CDM model, illustrating the application of a simple dynamical systems approach to a toy model, and providing a brief introduction to the Generalized SU(2) Proca theory. We demonstrated how even minor modifications to a model can significantly increase its complexity, necessitating advanced methods to derive meaningful insights and reconstruct the cosmic history. Chapter 2 further developed these concepts by focusing on an anisotropic Tachyon field, showcasing the use of numerical techniques within the dynamical systems framework to analyze its behavior and explore its implications for cosmic evolution.
@@ 3.2.3 Stability Analysis
As discussed in Section 1.5, the stability of fixed points can be determined by analyzing the eigenvalues of the Jacobian matrix, defined as: [^4] Alternative techniques, such as Lyapunov exponents [71] or Malkin’s criterion [136], may be required if linear stability analysis fails.
@@ 3.4 Conclusions
In this thesis, we employed the dynamical systems approach to explore two different cosmological models, each addressing distinct challenges in understanding the expansion history of the Universe. In Chapter 2, we developed a numerical framework for analyzing anisotropic dark energy models, specifically focusing on the interaction between a scalar tachyon field and a vector field in a Bianchi I background. This method, which does not rely on analytical fixed points, enabled the identification of regions within the parameter space where accelerated solutions could act as attractors of the system. The flexibility of this approach makes it applicable to any DE scenario, thereby demonstrating its general applicability.
In Chapter 3, we investigated the Generalized SU(2) Proca (GSU2P) theory, with particular focus on its potential to drive both primordial inflation and late-time cosmic acceleration. While the dynamical systems approach revealed regions where accelerated solutions could theoretically exist, it ultimately helped us rule out the GSU2P model as a viable candidate for a complete cosmological acceleration history. The model’s failure to produce stable attractors and maintain consistent cosmological evolution across different epochs prevented a smooth transition from inflation to a radiation-dominated phase, as well as from a matter-dominated phase to accelerated expansion. Thus, the dynamical systems approach served to discard this model as a viable candidate for cosmological history, highlighting its limitations in replicating the observed evolution of the Universe. In particular, the presence of singularities in phase space, such as those resulting from the regularization of dynamical equations, can disrupt the integrability and evolution of cosmological models. Depending on the behavior of the system during numerical integration, such singularities can halt its progression or introduce additional instabilities, ultimately limiting the model’s ability to replicate a consistent and observationally accurate cosmological history. These challenges underscore the need for rigorous examination and refinement of models to ensure their robustness across different epochs of cosmic evolution.